The magnetic moment of the Earth is approximately

8.00 × 1022 A · m2.

Imagine that the planetary magnetic field were caused by the complete magnetization of a huge iron deposit with density 7 900 kg/m3 and approximately

8.50 × 1028 iron atoms/m3.

(a) How many unpaired electrons, each with a magnetic moment of

9.27 × 10−24 A · m2,

would participate?


(b) At two unpaired electrons per iron atom, how many kilograms of iron would be present in the deposit?
Answer #1

No. of unpaired electrons = Total Magnetic moment/magnetic moment of each unpaired electron

No. = 8*10^22/(9.27*10^-24)

No = 8.63*10^45

Each iron atom has two unpaired electrons, so the number of iron atoms required is

= 8.63*10^45/2 = 4.315*10^45

B.

Mass = No. of participating iron atoms*density/Volume

Mass = 4.315*10^45*7900/(8.5*10^28)

Mass = 4.01*10^20 kg